3/x = 7/8
7x=3*8
7x=24
x=24/7=3 3/7

Found 2 solutions by checkley75, PaulBeckles:
Answer by checkley75(3666)

3/x = 7/8
7X=3*8
7X=24
X=24/7 OR 3 3/7

Answer by PaulBeckles(1)